# Extension property for partial automorphisms of the $n$-partite and semigeneric tournaments

### v3

@@ 1. Introduction
In 1992, Hrushovski [Hru92] established that the class of all finite graphs has EPPA. This result was used by Hodges, Hodkinson, Lascar, and Shelah to show the small index property for the random graph [HHLS93]. After this, the quest of studying EPPA continued with a series of papers including [Her95, Her98, HL00, HO03, Sol05, Ver08, ABWH+17, HKN18, Con19, HKN19, Kon19, Kon20, EHKN20, Ott20, EHN21, HKN22, BWCHK25].
@@ 1.1. Homogeneous structures
One can show that if a hereditary isomorphism-closed class $\mathcal{C}$ of finite relational structures has countably many members up to isomorphism, the joint embedding property (that is, for every $\mathbf{A},\mathbf{B}\in\mathcal{C}$ there is $\mathbf{C}\in\mathcal{C}$ which embeds both $\mathbf{A}$ and $\mathbf{B}$), and EPPA then it has the amalgamation property (that is, for every $\mathbf{A},\mathbf{B}_{1},\mathbf{B}_{2}\in\mathcal{C}$ and embeddings $\alpha_{i}\colon\mathbf{A}\to\mathbf{B}_{i}$ for $i\in\{1,2\}$, there is $\mathbf{C}\in\mathcal{C}$ with embeddings $\beta_{i}\colon\mathbf{B}_{i}\to\mathbf{C}$ for $i\in\{1,2\}$ such that $\beta_{1}\alpha_{1}=\beta_{2}\alpha_{2}$), see for example [HK25, Proposition 7.1].
@@
Semigeneric tournaments and $n$-partite tournaments appear in Cherlin’s classification of countable homogeneous directed graphs [Che98] and are one of the few known examples of amalgamation classes having EPPA where this property does not follow by a direct application of the Herwig–Lascar theorem [HL00], or more generally, its strengthening by Hubička, Konečný and Nešetřil [HKN22] (in addition to two-graphs [EHKN20], finite permutation groups [Sin17], and certain antipodal metrically homogeneous graphs [Kon20]). As discussed in [PS18], Theorems 1.1 and 1.2 imply that in order to fully classify which homogeneous directed graphs have EPPA, it now remains to decide EPPA for the class of all finite tournaments, the class of all finite directed graphs omitting an independent set of size $k$ ($k\geq 2$), and the class of all finite double covers of tournaments.
@@ 4.1. Background
As we have seen in the proof of Theorem 1.5, they do not have APA over the empty structure. The same situation happens for equivalences with $n$ equivalence classes and it is a consequence of these classes having non-trivial algebraic closure of the empty set after eliminating imaginaries (namely, the representatives of the equivalence classes are in the algebraic closure). If one only considers automorphisms which agree on $\mathrm{acl}(\emptyset)$ then it is possible to do the whole construction and obtain generic automorphisms with a fixed+ action on $\mathrm{acl}(\emptyset)$.
@@ 5.1. Witness construction
Let $\triangleleft$ be the order on part valuation functions defined by $g\triangleleft h$ if and only if $g\neq h$ and $g(v)=0$ for $v\in A$ being the $\leq$-least vertex for which $g(v)\neq h(v)$. Note that $h_{0}$ is the minimum of $\triangleleft$. Let $\prec$ be a linear order on $Q$ defined by $(A_{i},h)\prec(A_{j},h^{\prime})$ if and only if either $i<j$ or $i=j$ and $h\triangleleft h^{\prime}$.
@@ 7.1. Coherent EPPA
There is a strengthening of EPPA called coherent EPPA which was introduced by Siniora and Solecki [SS19] with the goal of generalizing the proof of Bhattacharjee and Macpherson that the automorphism group of the countable random graph contains a dense locally finite subgroup [BM05]: If $\mathbf{B}$ is an EPPA-witness for $\mathbf{A}$, we say that it is coherent if there is a map $\Psi$ from partial automorphisms of $\mathbf{A}$ to automorphisms of $\mathbf{B}$ such that $\Psi(f)$ extends $f$, and $\Psi(gf)=\Psi(g)\Psi(f)$ for every pair $f,g$ of partial automorphisms of $\mathbf{A}$ with $\mathop{\mathrm{Range}}\nolimits(f)=\mathop{\mathrm{Dom}}\nolimits(g)$. A class $\mathcal{C}$ has coherent EPPA if for every $\mathbf{A}\in\mathcal{C}$ there is a coherent EPPA-witness $\mathbf{B}\in\mathcal{C}$ for $\mathbf{A}$.

### v4

@@ 1. Introduction
In 1992, Hrushovski [Hru92] established that the class of all finite graphs has EPPA. This result was used by Hodges, Hodkinson, Lascar, and Shelah to show the small index property for the random graph [HHLS93]. After this, the quest of studying EPPA continued with a series of papers including [Her95, Her98, HL00, HO03, Sol05, Ver08, ABWH+26, HKN18, Con19, HKN19, Kon19, Kon20, EHKN20, Ott20, EHN21, HKN22, BWCHK25].
@@ 1.1. Homogeneous structures
One can show that if a hereditary isomorphism-closed class $\mathcal{C}$ of finite relational structures has countably many members up to isomorphism, the joint embedding property (that is, for every $\mathbf{A},\mathbf{B}\in\mathcal{C}$ there is $\mathbf{C}\in\mathcal{C}$ which embeds both $\mathbf{A}$ and $\mathbf{B}$), and EPPA then it has the amalgamation property (that is, for every $\mathbf{A},\mathbf{B}_{1},\mathbf{B}_{2}\in\mathcal{C}$ and embeddings $\alpha_{i}\colon\mathbf{A}\to\mathbf{B}_{i}$ for $i\in\{1,2\}$, there is $\mathbf{C}\in\mathcal{C}$ with embeddings $\beta_{i}\colon\mathbf{B}_{i}\to\mathbf{C}$ for $i\in\{1,2\}$ such that $\beta_{1}\alpha_{1}=\beta_{2}\alpha_{2}$), see for example [HK26, Proposition 7.1].
@@
Semigeneric tournaments and $n$-partite tournaments appear in Cherlin’s classification of countable homogeneous directed graphs [Che98] and are one of the few known examples of amalgamation classes having EPPA where this property does not follow by a direct application of the Herwig–Lascar theorem [HL00], or more generally, its strengthening by Hubička, Konečný and Nešetřil [HKN22] (in addition to two-graphs [EHKN20], finite permutation groups [Sin17], and certain antipodal metrically homogeneous graphs [Kon20]). As discussed in [PS20], Theorems 1.1 and 1.2 imply that in order to fully classify which homogeneous directed graphs have EPPA, it now remains to decide EPPA for the class of all finite tournaments, the class of all finite directed graphs omitting an independent set of size $k$ ($k\geq 2$), and the class of all finite double covers of tournaments.
@@ 4.1. Background
As we have seen in the proof of Theorem 1.5, they do not have APA over the empty structure. The same situation happens for equivalences with $n$ equivalence classes and it is a consequence of these classes having non-trivial algebraic closure of the empty set after eliminating imaginaries (namely, the representatives of the equivalence classes are in the algebraic closure). If one only considers automorphisms which agree on $\mathrm{acl}(\emptyset)$ then it is possible to do the whole construction and obtain generic automorphisms with a fixed action on $\mathrm{acl}(\emptyset)$.
@@ 5.1. Witness construction
Let $\triangleleft$ be the order on part valuation functions defined by $g\triangleleft h$ if and only if $g\neq h$ and $g(v)=0$ for $v\in A$ being the $\leq$-least vertex for which $g(v)\neq h(v)$. Note that $h_{0}$ is the minimum of $\triangleleft$. Let $\prec$ be the linear order on $Q$ defined by $(A_{i},h)\prec(A_{j},h^{\prime})$ if and only if either $i<j$ or $i=j$ and $h\triangleleft h^{\prime}$.
@@ 7.1. Coherent EPPA
There is a strengthening of EPPA called coherent EPPA which was introduced by Siniora and Solecki [SS20] with the goal of generalizing the proof of Bhattacharjee and Macpherson that the automorphism group of the countable random graph contains a dense locally finite subgroup [BM05]: If $\mathbf{B}$ is an EPPA-witness for $\mathbf{A}$, we say that it is coherent if there is a map $\Psi$ from partial automorphisms of $\mathbf{A}$ to automorphisms of $\mathbf{B}$ such that $\Psi(f)$ extends $f$, and $\Psi(gf)=\Psi(g)\Psi(f)$ for every pair $f,g$ of partial automorphisms of $\mathbf{A}$ with $\mathop{\mathrm{Range}}\nolimits(f)=\mathop{\mathrm{Dom}}\nolimits(g)$. A class $\mathcal{C}$ has coherent EPPA if for every $\mathbf{A}\in\mathcal{C}$ there is a coherent EPPA-witness $\mathbf{B}\in\mathcal{C}$ for $\mathbf{A}$.